# On the Parameter Identifiability of Partially Observed Linear Causal Models

### v1

@@ 3.2 Graphical Condition for Structure Identifiability
Let $\mathbf{V}\in\mathbf{V}_{\mathcal{G}}$ be a set of variables, where $l$ out of $|\mathbf{V}|$ are latent, and the remaining $|\mathbf{V}|-l$ are observed. $\mathbf{V}$ is an atomic cover if $\mathbf{V}$ contains a single observed variable, or if the following conditions hold:
@@
    There exists a set of covers $\mathcal{N}$, with $||\mathcal{N}||\geq l+1$, such that every element in $\cup_{\mathbf{N}\in\mathcal{N}}\mathbf{N}$ is a neighbour of $\mathbf{V}$ and $(\cup_{\mathbf{N}\in\mathcal{N}}\mathbf{N})\cap(\cup_{\mathbf{C}\in\mathcal{C}}\mathbf{C})=\emptyset$.
@@
    There is not a partition of $\mathbf{V}=\mathbf{V_{1}}\cup\mathbf{V_{2}}$ such that both $\mathbf{V_{1}}$ and $\mathbf{V_{2}}$ are atomic covers.
Figure 3: An illustrative graph that satisfies the conditions for structure-identifiability. At the same time, it also satisfies the condition for parameter identifiability - given the structure and $\Sigma_{\mathbf{X}_{\mathcal{G}}}$, all the parameters are identifiable only upto group sign indeterminacy.
@@
Consider the graph in Fig. 3. $\mathbf{V}=\{\mathsf{L_{1}},\mathsf{X_{6}}\}$ is an atomic cover. This is because there exist $\mathcal{C}=\{\{\mathsf{X_{9}}\},\{\mathsf{X_{10}}\}\}$ with $||\mathcal{C}||\geq l+1=2$ such that (i) in Def. 5 is satisfied. And there exist $\mathcal{N}=\{\{\mathsf{X_{11}}\},\{\mathsf{X_{12}}\}\}$ (or, $\mathcal{N}=\{\{\mathsf{X_{4}}\},\{\mathsf{X_{5}}\}\}$) with $||\mathcal{N}||\geq l+1=2$ such that (ii) in Def. 5 is satisfied. We can also find that (iii) in Def. 5 is satisfied. Thus $\{\mathsf{L_{1}},\mathsf{X_{6}}\}$ is an atomic cover. Another example would be in Figure 8, where $\{\mathsf{L_{1}},\mathsf{L_{2}}\}$ is an atomic cover.
@@
Consider Figure 3. All latent variables in the graph belong to at least one atomic cover and thus Condition 1 is satisfied. Plus, Condition 2 is also satisfied. This is because the sets of variables $\mathbf{V}$, $\mathbf{V_{1}}$, $\mathbf{V_{2}}$, and $\mathbf{T}$ that satisfy (i) and (ii) in Condition 2 are $\mathbf{V}=\{\mathsf{X_{12}}\}$, $\mathbf{V_{1}}=\{\mathsf{L_{1}},\mathsf{X_{6}}\}$, $\mathbf{V_{2}}=\{\mathsf{L_{2}}\}$, and $\mathbf{A}=\{\mathsf{X_{4}},\mathsf{X_{5}}\}$, and we also have $|\mathbf{V}|+|\mathbf{A}|\geq|\mathbf{V_{1}}|+|\mathbf{V_{2}}|$. Therefore, the graph in Figure 3 satisfies both Conditions 1 and 2.
The identifiability theory of structure is as follows. For a graph $\mathcal{G}$, if Condition 1 and Condition 2 are satisfied, then asymptotically the structure is identifiable upto the Markov equivalence class (MEC) of $\mathcal{O}_{\text{min}}(\mathcal{O}_{s}(\mathcal{G}))$ (definitions of $\mathcal{O}_{\text{min}}(\cdot)$ and $\mathcal{O}_{s}(\cdot)$ can be found in Appendix B.3). Roughly speaking, the underlying causal structure of $\mathcal{G}$ can be identified except that the directions of some edges cannot be determined. Next, we will show that, given any DAG in the identified equivalence class together with $\Sigma_{\mathbf{X}_{\mathcal{G}}}$, the parameters of the model are also identifiable, if certain conditions are satisfied.
@@ 3.3 Identifiability of Parameters
Theorem 1 (Sufficient Condition for Parameter identifiability (upto group sign)).
Assume that $\mathcal{G}$ satisfies Conditions 1 and 2 and thus the structure can be identified upto the MEC of $\mathcal{O}_{\text{min}}(\mathcal{O}_{s}(\mathcal{G}))$. For any DAG in the equivalence class, the parameters are identifiable, if both the following hold:
@@
    If an atomic cover $\mathbf{V}=\mathbf{X}\cup\mathbf{L}$ satisfies $|\mathbf{L}|\neq 0$ and $|\mathbf{X}|\geq 1$, then there exists $\mathbf{\hat{X}}\subseteq\mathbf{X}_{\mathcal{G}}$ such that $\mathbf{\hat{X}}$ d-separates $\mathbf{X}$ and $\mathbf{L}$.
Theorem 1 provides a sufficient condition such that the parameters are identifiable. We note that this sufficient condition is pretty close to a necessary and sufficient condition, and we will detail this point later after we introduced some necessary conditions that can guide us under which scenarios parameters are guaranteed to be not identifiable. Now, for a better understanding of Theorem 1, we provide an example of it as follows.
@@
The graph $\mathcal{G}$ in Figure 3 satisfies the conditions for parameter identifiability in Theorem 1. Specifically, condition (i) in Theorem 1, is satisfied as all atomic covers contain no more than one latent variable. Plus, condition (ii) in Theorem 1 is also satisfied, as the atomic cover $\mathbf{V}=\mathbf{X}\cup\mathbf{L}=\{\mathsf{L_{1}}\}\cup\{\mathsf{X_{6}}\}$ satisfies $|\mathbf{L}|\neq 0$ and $|\mathbf{X}|\geq 1$ and there exist $\mathbf{\hat{X}}=\{\mathsf{X_{4}},\mathsf{X_{5}}\}$ such that $\mathbf{\hat{X}}$ d-sep $\mathbf{X}$ and $\mathbf{L}$. Therefore, the parameters are identifiable for the graph in Figure 3.
Next, we discuss under which conditions the parameters are guaranteed to be not identifiable. As discussed in Section 3.1, there are three kinds of indeterminacy. The first one can be solved by assuming unit variance of latent variables while the second one group sign indeterminacy is rather trivial such that we still consider parameters as identifiable even if group sign indeterminacy exists. Therefore, we will focus on the third one, orthogonal transformation indeterminacy, in what follows.
@@
Example 5 (Example for Thm. 2).
Consider Fig. 8. The graph satisfies the conditions in Thm. 2 as the parents and children of $\mathsf{L_{1}}$ and $\mathsf{L_{2}}$ are exactly the same. Therefore, there must exist orthogonal transformation indeterminacy for the edge coefficients $F$ and thus the parameters are not identifiable.
@@
Assume $\mathcal{G}$ satisfies Conditions 1 and 2 and thus the structure can be identified upto the MEC of $\mathcal{O}_{\text{min}}(\mathcal{O}_{s}(\mathcal{G}))$. For any DAG $\mathcal{G}$ in the equivalence class, for $\mathcal{G}$’s parameters to be identifiable, every atomic cover must contain no more than one latent variable.
Now we are ready to discuss the necessity of the graphical condition proposed in Theorem 1.
@@
Theorem 1 provides a sufficient condition for parameter identifiability and it is pretty close to a necessary and sufficient one with reasons as follows. First, condition (i) in Theorem  1 is provably necessary: by Corollary 2, for parameters to be identifiable, one has to assume (i) in Theorem  1. Second, if condition (ii) in Theorem  1 does not hold, then there are only some rare cases where parameters can be identified.
We would like to mention that establishing a necessary and sufficient condition is always highly non-trivial in various tasks. For example, for the identification of linear non-Gaussian causal structure with latent variables, researchers initially developed sufficient conditions with three pure children in [44], later relaxed to two in [10, 56], before ultimately achieving both necessary and sufficient conditions in [1]. Similarly, for parameter identification, although the condition we proposed is not a necessary and sufficient one, it has considerable extents of necessity, and could be expected to serve as a stepping stone towards tighter and ultimately the necessary and sufficient condition for the field.
Identifiability theory often focuses on the asymptotic case, i.e., we assume that we know the structure and the population covariance matrix $\Sigma_{\mathbf{X}_{\mathcal{G}}}$. However, in practice, we only have access to i.i.d. data with finite sample size and thus only have the sample covariance matrix. Therefore, in the next section, we will propose a novel method to estimate the parameters in the finite sample cases.
@@ 4.1 Objective
Our goal is to estimate $F$ in Definition 1, given the causal structure $\mathcal{G}$ and observational data. The key is to parameterize the population covariance $\Sigma_{\mathbf{X_{\mathcal{G}}}}$ using $\theta=(F,\Omega)$ and then maximize the likelihood of observed sample covariance $\hat{\Sigma}_{\mathbf{X_{\mathcal{G}}}}$. To make this technically precise, we provide a closed-form expression of $\Sigma_{\mathbf{X}}$ in terms of $\theta$ in the following proposition, with a proof given in Section A.5.
Proposition 1 (Parameterization of Population Covariance Matrix).
@@
Consider the model defined in Definition 1. Let $M\coloneqq\left(I-A-B(I-D)^{-1}C\right)^{-1}$ and $N\coloneqq\left((I-A)C^{-1}(I-D)-B\right)^{-1}$. Then, the population covariance matrices of $\mathbf{L}_{\mathcal{G}}$ and $\mathbf{X}_{\mathcal{G}}$ can be formulated as
$\Sigma_{\mathbf{L}_{\mathcal{G}}}=$ $M^{T}\Omega_{\mathbf{L}_{\mathcal{G}}}M+N^{T}\Omega_{\mathbf{X}_{\mathcal{G}}}N,$ (1)
$\Sigma_{\mathbf{X}_{\mathcal{G}}}=$ $(I-D)^{-T}\Big{(}\Omega_{\mathbf{X}_{\mathcal{G}}}+B^{T}\Sigma_{\mathbf{L}_{\mathcal{G}}}B+\Omega_{\mathbf{X}_{\mathcal{G}}}NB+B^{T}N^{T}\Omega_{\mathbf{X}_{\mathcal{G}}}\Big{)}(I-D)^{-1}.$ (2)
@@
The formulations of $\Sigma_{\mathbf{L}_{\mathcal{G}}}$ and $\Sigma_{\mathbf{X}_{\mathcal{G}}}$ are rather complicated due to the general scenario we considered, i.e., latent variables can be the cause or the effect of latent and observed variables. That is, the submatrices $A,B,C$ and $D$ defined in the above proposition can all have nonzero entries. In most existing works, at least one of these submatrices are assumed to be zero. For instance, factor analysis assumes that $A,C$ and $D$ are zero, while [31] assumes that $A$ and $C$ are zero. Furthermore, Proposition 1 also provides insight into the indeterminacy involved when identifying the parameters, such as the indeterminacy of the variance of latent variables in Lemma 1 and the rotational indeterminacy in Theorem 2.
@@
Similar to the typical factor analysis setting [43, 7, 20], we assume that $\epsilon_{\mathbf{V_{\mathcal{G}}}}$ are all gaussian and thus $\mathbf{X}_{\mathcal{G}}$ are jointly Gaussian. With the parametric form of the likelihood, we find the causal coefficients $F$ by minimizing the negative log-likelihood of observational data:
$\min_{F,\Omega}~{}(K/2)(\operatorname{tr}((\Sigma_{\mathbf{X}_{\mathcal{G}}})^{-1}\hat{\Sigma}_{\mathbf{X}_{\mathcal{G}}})+\log\det\Sigma_{\mathbf{X}_{\mathcal{G}}})\quad\text{subject to}~{}(\Sigma_{\mathbf{L}_{\mathcal{G}}})_{ii}=1,~{}i\in[m],$ (3)
where $K$ is the number of i.i.d. observations. Note that the entries of matrix $F$ that do not correspond to an edge in $\mathcal{G}$ are constrained to be zero during the optimization. In our empirical studies we employ gradient descent for the optimization owing to its simplicity and rely on multiple random initializations to address nonconvexity. Note that the constraint in Equation 3 that all latent variables have unit variance is crucial to deal with the variance indeterminacy of latent variables, and yet, it may not be straightforward to realize this constraint. One classical solution is to use the Lagrange multiplier technique. However, adding this term will make the nonconvex optimization problem even harder (which is empirically validated in our experiments in Section 5).
It is desirable to parameterize the population covariance such that the constraint in Equation 3 can be automatically satisfied. This spirit follows from factor analysis, where the population covariance is $\Sigma_{\mathbf{X}_{\mathcal{G}}}=B^{T}\Sigma_{\mathbf{L}_{\mathcal{G}}}B+\Omega_{\mathbf{X}_{\mathcal{G}}}$, and latent variables are independent; thus, the unit variance constraint of latent variables can be satisfied by simply taking $\Sigma_{\mathbf{L}_{\mathcal{G}}}=I$. However, in our setting, latent variables are not independent and the parameterization form of $\Sigma_{\mathbf{X}_{\mathcal{G}}}$ in Proposition 1 is complicated, both of which makes it not straightforward to realize the same goal. To this end, we introduce the following way to parameterize $\Sigma_{\mathbf{X}_{\mathcal{G}}}$ using $F$, such that the required constraint is automatically satisfied.

### v2

@@ 3.2 Graphical Condition for Structure Identifiability
Let $\mathbf{V}\in\mathbf{V}_{\mathcal{G}}$ be a set of variables, where $l$ out of $|\mathbf{V}|$ are latent, and the remaining $|\mathbf{V}|-l$ are observed. $\mathbf{V}$ is an atomic cover if $\mathbf{V}$ is a single observed variable, or if the following conditions hold:
@@
    There exists a set of covers $\mathcal{N}$ with $||\mathcal{N}||\geq l+1$, s.t. $(\cup_{\mathbf{N}\in\mathcal{N}}\mathbf{N})\cap(\cup_{\mathbf{C}\in\mathcal{C}}\mathbf{C})=\emptyset$, every element in $\cup_{\mathbf{N}\in\mathcal{N}}\mathbf{N}$ is a neighbour of every element in $\mathbf{V}$, and $\mathbf{V}$ d-separates $\mathcal{N}$ and $\mathcal{C}$.
@@
    There does not exist a partition $\mathcal{P}$ of $\mathbf{V}$, s.t., all elements in $\mathcal{P}$ are atomic covers.
Figure 3: An illustrative graph that satisfies the conditions for structure-identifiability. At the same time, it also satisfies the condition for parameter identifiability - given the structure and $\Sigma_{\mathbf{X}}$, all the parameters are identifiable only up to group sign indeterminacy.
@@
Consider the graph in Fig. 3. $\mathbf{V}=\{\mathsf{L_{1}},\mathsf{X_{6}}\}$ is an atomic cover. This is because there exist $\mathcal{C}=\{\{\mathsf{X_{9}}\},\{\mathsf{X_{10}}\}\}$ with $||\mathcal{C}||\geq l+1=2$ such that (i) in Def. 4 is satisfied. And there exist $\mathcal{N}=\{\{\mathsf{X_{11}}\},\{\mathsf{X_{12}}\}\}$ (or, $\mathcal{N}=\{\{\mathsf{X_{4}}\},\{\mathsf{X_{5}}\}\}$) with $||\mathcal{N}||\geq l+1=2$ such that (ii) in Def. 4 is satisfied. We can also find that (iii) in Def. 4 is satisfied. Thus $\{\mathsf{L_{1}},\mathsf{X_{6}}\}$ is an atomic cover. Another example would be in Figure 8, where $\{\mathsf{L_{1}},\mathsf{L_{2}}\}$ is an atomic cover.
@@
Consider Figure 3. All latent variables in the graph belong to at least one atomic cover and thus Condition 1 is satisfied. Plus, Condition 2 is also satisfied. This is because the sets of variables $\mathbf{V}$, $\mathbf{V_{1}}$, $\mathbf{V_{2}}$, and $\mathbf{T}$ that satisfy (i) and (ii) in Condition 2 are $\mathbf{V}=\{\mathsf{X_{12}}\}$, $\mathbf{V_{1}}=\{\mathsf{L_{1}},\mathsf{X_{6}}\}$, $\mathbf{V_{2}}=\{\mathsf{L_{2}}\}$, and $\mathbf{T}=\{\mathsf{X_{4}},\mathsf{X_{5}}\}$, and we also have $|\mathbf{V}|+|\mathbf{T}|\geq|\mathbf{V_{1}}|+|\mathbf{V_{2}}|$. Therefore, the graph in Figure 3 satisfies both Conditions 1 and 2.
The identifiability theory of structure is as follows. For a graph $\mathcal{G}$, if Condition 1 and Condition 2 are satisfied, then asymptotically the structure is identifiable up to the Markov equivalence class (MEC) of $\mathcal{O}_{\text{a}}(\mathcal{O}_{s}(\mathcal{G}))$ (definitions of $\mathcal{O}_{\text{a}}(\cdot)$ and $\mathcal{O}_{s}(\cdot)$ can be found in Appendix B.4). Roughly speaking, the underlying causal structure of $\mathcal{G}$ can be identified except that the directions of some edges cannot be determined. Next, we will show that, given any DAG in the identified equivalence class together with $\Sigma_{\mathbf{X}}$, the parameters of the model are also identifiable, if certain conditions are satisfied.
@@ 3.3 Identifiability of Parameters
Theorem 3 (Sufficient Condition for Parameter Identifiability (up to group sign), Based on Structure Identifiability).
Assume that $\mathcal{G}$ satisfies Conditions 1 and 2 and thus the structure can be identified up to the MEC of $\mathcal{O}_{\text{a}}(\mathcal{O}_{s}(\mathcal{G}))$. For any DAG in the equivalence class, the parameters are identifiable, if both the following hold:
@@
    If an atomic cover $\mathbf{V}=\mathbf{X}\cup\mathbf{L}$ satisfies $|\mathbf{L}|\neq 0$ and $|\mathbf{X}|\geq 1$, then all simple treks (Def. 5) between $\mathbf{L}$ and $\mathbf{X}$ do not contain any latent variables that are not in $\mathbf{L}$.
Theorem 3 provides a sufficient condition such that the parameters are identifiable. Now, for a better understanding of Theorem 3, we provide an example of it as follows.
@@
The graph $\mathcal{G}$ in Figure 3 satisfies the conditions for parameter identifiability in Theorem 3. Specifically, condition (i) in Theorem 3, is satisfied as all atomic covers contain no more than one latent variable. Plus, condition (ii) in Theorem 3 is also satisfied, as the atomic cover $\mathbf{V}=\mathbf{X}\cup\mathbf{L}=\{\mathsf{L_{1}}\}\cup\{\mathsf{X_{6}}\}$ satisfies $|\mathbf{L}|\neq 0$ and $|\mathbf{X}|\geq 1$ and all simple treks between $\{\mathsf{L_{1}}\}$ and $\{\mathsf{X_{6}}\}$ contain only observed variables except $\{\mathsf{L_{1}}\}$. Therefore, the parameters are identifiable for the graph in Figure 3.
Next, we discuss under which conditions the parameters are guaranteed to be not identifiable. As discussed in Section 3.1, there are three kinds of indeterminacy. The first one can be solved by assuming unit variance of the noise terms of latent variables while the second one group sign indeterminacy is rather trivial such that we still consider parameters as identifiable even if group sign indeterminacy exists. Therefore, we will focus on the third one, orthogonal transformation indeterminacy, in what follows.
@@
Example 5 (Example for Thm. 4).
Consider Fig. 8. The graph satisfies the conditions in Thm. 4 as the parents and children of $\mathsf{L_{1}}$ and $\mathsf{L_{2}}$ are exactly the same. Therefore, there must exist orthogonal transformation indeterminacy for the edge coefficients $F$ and thus the parameters are not identifiable.
@@
Assume $\mathcal{G}$ satisfies Conditions 1 and 2 and thus the structure can be identified up to the MEC of $\mathcal{O}_{\text{a}}(\mathcal{O}_{s}(\mathcal{G}))$. For any DAG $\mathcal{G}$ in the equivalence class, for $\mathcal{G}$’s parameters to be identifiable, every atomic cover must contain no more than one latent variable.
@@
Condition (i) in Theorem  3 is provably necessary: by Corollary 2, for parameters to be identifiable, one has to assume (i) in Theorem  3.
Establishing a necessary and sufficient condition is always highly non-trivial in various tasks. For example, for the identification of linear non-Gaussian causal structure with latent variables, researchers initially developed sufficient conditions with three pure children in [46], later relaxed to two in [11, 58], before ultimately achieving both necessary and sufficient conditions in [1]. Similarly, for parameter identification, although the condition we proposed is not a necessary and sufficient one, it could serve as a stepping stone towards tighter and ultimately the necessary and sufficient condition for the field.
Below, we also provide a sufficient condition for parameter identifiability that does not rely on structure identifiability in Theorem 5. It is particularly useful when the structure is directly given by some domain experts.
Theorem 5 (Sufficient Condition for Parameter Identifiability (up to group sign) without Requiring Structure Identifiability).
In $\mathcal{G}$, if for every latent variable $\mathsf{L}$ there always exist another three distinct variables (which can be latent or observed), such that two of the three are pure children of $\mathsf{L}$ and the rest one is a neighbor of $\mathsf{L}$, then the parameters are identifiable.
Identifiability theory often focuses on the asymptotic case, i.e., we assume that we know the structure and the population covariance matrix $\Sigma_{\mathbf{X}}$. However, in practice, we only have access to i.i.d. data with finite sample size and thus only have the sample covariance matrix. Therefore, in the next section, we will propose a novel method to estimate the parameters in the finite sample cases.
@@ 4.1 Objective
Our goal is to estimate $F$ in Definition 1, given the causal structure $\mathcal{G}$ and observational data. The key is to parameterize the population covariance $\Sigma_{\mathbf{X}}$ using $\theta=(F,\Omega)$ and then maximize the likelihood of observed sample covariance $\hat{\Sigma}_{\mathbf{X}}$. To make this technically precise, we provide a closed-form expression of $\Sigma_{\mathbf{X}}$ in terms of $\theta$ in the following proposition, with a proof given in Section A.7.
Proposition 1 (Parameterization of Population Covariance).
@@
Consider the model defined in Def. 1. Let $M\coloneqq\left(I-F_{\mathbf{L}\mathbf{L}}-F_{\mathbf{L}\mathbf{X}}(I-F_{\mathbf{X}\mathbf{X}})^{-1}F_{\mathbf{X}\mathbf{L}}\right)^{-1}$ and $N\coloneqq\left((I-F_{\mathbf{L}\mathbf{L}})F_{\mathbf{X}\mathbf{L}}^{-1}(I-F_{\mathbf{X}\mathbf{X}})-F_{\mathbf{L}\mathbf{X}}\right)^{-1}$. Then, the population covariance matrices of $\mathbf{L}$ and $\mathbf{X}$ can be formulated as
$\Sigma_{\mathbf{L}}=$ $M^{T}\Omega_{\epsilon_{\mathbf{L}}}M+N^{T}\Omega_{\epsilon_{\mathbf{X}}}N,$ (1)
$\Sigma_{\mathbf{X}}=$ $(I-F_{\mathbf{X}\mathbf{X}})^{-T}\Big{(}F_{\mathbf{L}\mathbf{X}}^{T}\Sigma_{\mathbf{L}_{\mathcal{G}}}F_{\mathbf{L}\mathbf{X}}+\Omega_{\epsilon_{\mathbf{X}}}+\Omega_{\epsilon_{\mathbf{X}}}NF_{\mathbf{L}\mathbf{X}}+F_{\mathbf{L}\mathbf{X}}^{T}N^{T}\Omega_{\epsilon_{\mathbf{X}}}\Big{)}(I-F_{\mathbf{X}\mathbf{X}})^{-1}.$ (2)
@@
The formulations of $\Sigma_{\mathbf{L}}$ and $\Sigma_{\mathbf{X}}$ are rather complicated due to the general scenario we considered, i.e., latent variables can be the cause or the effect of latent and observed variables. That is, the submatrices $F_{\mathbf{L}\mathbf{L}},F_{\mathbf{L}\mathbf{X}},F_{\mathbf{X}\mathbf{L}}$ and $F_{\mathbf{X}\mathbf{X}}$ defined in the above proposition can all have nonzero entries. In most existing works, at least one of these submatrices are assumed to be zero. For instance, factor analysis assumes that $F_{\mathbf{L}\mathbf{L}},F_{\mathbf{X}\mathbf{L}}$ and $F_{\mathbf{X}\mathbf{X}}$ are zero, while [32] assumes that $F_{\mathbf{L}\mathbf{L}}$ and $F_{\mathbf{X}\mathbf{L}}$ are zero. Furthermore, Proposition 1 also provides insight into the indeterminacy involved when identifying the parameters, such as the indeterminacy of variance in Theorem 1 and the orthogonal transformation indeterminacy in Theorem 4.
@@
Similar to factor analysis [45, 7, 21], we assume $\epsilon_{\mathbf{V}}$ are Gaussian and thus $\mathbf{X}$ are jointly Gaussian. Thus, the negative log-likelihood of observational data can be formulated as
$\mathcal{L}=(K/2)(\operatorname{tr}((\Sigma_{\mathbf{X}})^{-1}\hat{\Sigma}_{\mathbf{X}})+\log\det\Sigma_{\mathbf{X}}),$ (3)
where $K$ is the number of i.i.d. observations. With the parameterized negative log-likelihood, we estimate the edge coefficients by minimizing the negative log-likelihood, as
$\hat{F},\hat{\Omega}=\arg\min_{F,\Omega}~{}\mathcal{L},\quad\text{subject to}~{}\Omega_{\epsilon_{\mathbf{L}}}=I,$ (4)
where the entries of matrix $F$ that do not correspond to an edge in $\mathcal{G}$ are constrained to be zero during the optimization.
Note that in Eq. (4) the constraint that the noise terms of latent variables have unit variance is crucial to deal with the variance indeterminacy defined in Theorem 1. In practice, it is also favorable to use another constraint to address the variance indeterminacy, i.e., the constraint that all the latent variables have unit variance. This leads to an alternative objective as
$\hat{F},\hat{\Omega}=\arg\min_{F,\Omega}~{}\mathcal{L},\quad\text{subject to}~{}(\Sigma_{\mathbf{L}})_{ii}=1,~{}i\in[m],$ (5)
where the entries of $F$ that do not correspond to an edge in $\mathcal{G}$ are also constrained to be zero.
Both objectives in Eqs. (4) and (5) can be employed, and yet using the second one gives rise to edge coefficients that are easier to understand. To be concerete, if we normalize all observed variables to have unit variance, then using Eq. (5) would give rise to $\hat{F}$ such that $-1\leq\hat{F}_{i,j}\leq 1,\forall i,j\in[m]$. An example can be found in Figure 4. However, it may not be straightforward to realize the constraint in Eq. (5). To this end, in the next section we introduce a way to parameterize $\Sigma_{\mathbf{X}}$ using $F$, such that the required constraint in Eq. (5) can be automatically satisfied. Later in Section 5.2, we also empirically compare the performance of using Eq. (4) with that of using Eq. (5).